# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 8.1. The Fredholm and invertibility properties of $\boldsymbol{\Xi}$
$\color[rgb]{1,0,1}\widehat{V}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V\succ 0\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V_{0}\succ 0\quad\mbox{on }M\,.$
Notice that we are not assuming that, globally, $V$ and $V_{0}$ are non-negative. This will be done in Assumption 8.3, which is, hence slightly stronger than Assumption 8.1. We also recall that, if $a$ is a multiplication operator, then we use also the notation $\color[rgb]{1,0,1}\widehat{a}$ for ${\mathcal{R}_{\infty}}(a)$ (see Equation (5.6) for a justification of this notation). We also recall (see Definition 2.13) that the explicit forms of the last two assumptions amount to the following:
@@
    $\color[rgb]{1,0,1}\widehat{V}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}(x)$ is invertible (i.e., $V_{0}(x)>0$ in $M_{n}(\mathbb{C})$).
@@
    $\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}_{0}(x)$ is invertible (i.e., $V_{0}(x)\in(0,\infty)$).
@@ 8.2. The Moore-Penrose pseudoinverse of $\boldsymbol{\Xi}$
Let $\mathcal{N}$ be the kernel of $\boldsymbol{\Xi}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$. The elliptic regularity of $\boldsymbol{\Xi}$ implies that $\mathcal{N}$ will consist of smooth sections. Assume that $\mathcal{N}\subset L^{2}(M;TM\oplus\mathbb{C})$ and that it is finite dimensional and let $p_{\mathcal{N}}$ be the $L^{2}$–orthogonal projection onto $\mathcal{N}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$ in Definition 6.11. Assume also that there exists a pseudodifferential operator
@@
Let $M$ be a manifold with straight cylindrical ends and assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ (Assumption 8.1). Then the following properties hold.
@@
Assumption 8.1 implies that ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ satisfies the assumptions of Theorem 7.14, which then yields right away the first point. The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is elliptic by Proposition 6.13. Since ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ is invertible, Theorem 6.15 implies that $\boldsymbol{\Xi}$ is Fredholm. The last point follows from the Beals-type result proved in [67, Theorem 7.4] (this result can also be found in [63]). ∎
@@
Let $U\in H^{1}(M;TM)\oplus L^{2}(M)$ be such that $\boldsymbol{\Xi}U=0$ on $M$. Then $U\in H^{2}(M;TM)\oplus H^{1}(M)$, by elliptic regularity. Let us notice that the fact that $V$ and $V_{0}$ satisfy the non-vanishing assumption on $M$ (Assumption 8.3) and that $M$ is connected implies that $V_{0}\not\equiv 0$ and $V\succ 0$ on $M$. The assumption $\boldsymbol{\Xi}U=0$ then allows us to use Corollary 2.15 to conclude that $U=0$. Thus $\boldsymbol{\Xi}$ is invertible between the indicated spaces, and hence also as a pseudodifferential operator. The fact that its inverse belongs to the set $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ follows from a “Beals-type theorem” (Theorem 7.4 in [67]). The last stated isomorphism follows from the mapping properties of operators in the spaces $\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM)$ and in $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM)$, as in the last corollaries. This completes the proof. ∎
The invertibility of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ in the case $M$ closed is obtained when both $V$ and $V_{0}$ do not vanish identically on $M$. Moreover, the condition that $V$ does not vanish identically on $M$ can be replaced with the condition that $M$ does not have non-vanishing Killing vector fields. These results were proved in [69, Theorem 6.1(4)]. In relation to this result, the reader might have noticed that we are not considering the case when $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ has straight cylindrical ends but does not have non-vanishing Killing vector fields. In fact, this condition is not very useful for manifolds with straight cylindrical ends because, even if $M$ does not have non-vanishing Killing vector fields, its “end” $M^{\prime}\times\mathbb{R}$ will always have non-vanishing Killing vector fields, for instance $\partial_{t}$ (the infinitesimal generator of the translations along $\mathbb{R}$). (See Subsection 12.2 for more on the case $M$ closed.)
@@ 8.4. Jump relations
For the Stokes operator $\boldsymbol{\Xi}_{0,0}$, see also[39], [62, Lemma 3.1], [70, (6.1), (6.2)], [99, Theorem 3.1], [100], [134, Lemma 1.3]. Note that $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\frac{1}{2}+\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$. Therefore $\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\boldsymbol{S}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$. In addition, we obtain the following formula.
@@ 9. Calculations on a cylinder
We shall need definitions and results for the limit operators ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ that are analogous to those in the previous two sections. To this end, in this section we consider the particular case when our manifold $M$ with straight cylindrical ends is actually a cylinder $M=\mathfrak{S}\times\mathbb{R}$ with a product metric. We do not assume $\mathfrak{S}$ to be connected, though. Then, as in Section 7, we let $\Omega=\Omega^{\prime}\times\mathbb{R}\subset M$ be a product type open subset, and hence with straight cylindrical ends, where $\Omega^{\prime}\subset\mathfrak{S}$. Let $\Gamma^{\prime}:=\partial\Omega^{\prime}$. In this section, in addition to $V$ and $V\geq 0$ smooth, we assume that they are translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$. Also, in this section, we use Assumption 8.1, that is that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (that is, they have points where they are strictly positive in every connected component of the space of definition, see Definition 2.13). Also, recall that $\Omega$ is on one side of its boundary (Assumption 2.7), and hence $\Omega^{\prime}$ is also on one side of its boundary.
@@ 9.1. Layer potentials and jump relations for the indicial operators
Assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (Assumption 8.1). Let $\boldsymbol{h}\in H^{s}(\Gamma^{\prime};TM)$, $s\in\mathbb{R}$. The indicial layer potentials operators associated to $\widehat{\boldsymbol{\Xi}}(\tau)$ are defined by
@@
Let $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$, $\Gamma^{\prime}:=\partial\Omega^{\prime}$, $\color[rgb]{1,0,1}\widehat{\mathfrak{f}}:=\frac{\color[rgb]{1,0,1}\widehat{V}_{0}+1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$ $\color[rgb]{1,0,1}\widehat{\mathfrak{g}}:=\frac{1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$. Let $\boldsymbol{K}:=\boldsymbol{P}_{0}\in\Psi^{0}(\Gamma^{\prime};TM)$ be as in Equation (8.6). Then
@@
    $\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\,=\,\widehat{\mathcal{A}}_{0}(\tau)\boldsymbol{h}^{\prime}$, and $\widehat{\mathcal{A}}_{0}(\tau)$ is elliptic with self-adjoint symbol $\sigma_{-1}(\mathcal{A}_{0}(\tau);\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$.
@@
    $\mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}=\big(\mp\frac{\color[rgb]{1,0,1}\widehat{\mathfrak{g}}}{2}\boldsymbol{\nu}^{\sharp}+\widehat{\mathcal{C}}_{0}(\tau)\big)\boldsymbol{h}^{\prime}$.
@@ 9.2. Invertibility of $\boldsymbol{S}$ on a cylinder
  • $V$ and $V\geq 0$ are smooth, translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
  • $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1, that is, $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$).
@@
    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and
@@
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
@@
Let us show now that $\ker\widehat{\boldsymbol{S}\,}(0)=0$. Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Our assumptions imply that $\widehat{\boldsymbol{\Xi}}(0)$ is invertible (Proposition 7.11), and hence the indicial single layer potential $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{0}(\boldsymbol{h}^{\prime})$ is defined (see Definition 9.2). Since $\widehat{\boldsymbol{S}\,}(0)$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity. Proposition 9.5 for $s=1/2$ then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}{\,}^{\prime}$ satisfy
@@
We also know that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{\pm}^{\prime}$ by Proposition 8.10. Proposition 9.6 gives then that $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Therefore $\boldsymbol{u}=0$ in $\Omega_{-}^{\prime}$ and in $\Omega_{+}^{\prime}$, by Corollary 7.7. The same corollary gives that $p$ is constant on each connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. Because $\color[rgb]{1,0,1}\widehat{V}_{0}$ is not identically equal to zero on any connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$, we obtain that, in fact, $p=0$. We have thus obtained that $U=0$ and hence $\widehat{\boldsymbol{S}\,}(0)$ is injective.
@@
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $V\succ 0$ and $V_{0}\succ 0$ on $M$ (equivalently, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$).
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $V$ and $V_{0}$ are smooth, translation invariant, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The proof is exactly the same as that of Proposition 9.7, except that we are using Corollary 7.9 (which only requires $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ to be non-negative) instead of Corollary 7.7. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. Then $\boldsymbol{S}$ is invertible, $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, and hence $\boldsymbol{S}:H^{s}(\Gamma;TM)\to H^{s+1}(\Gamma;TM)$ is an isomorphism for all $s\in\mathbb{R}$.

### 28 May 2026 v2

@@ 8.1. The Fredholm and invertibility properties of $\boldsymbol{\Xi}$
$\widehat{V}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V\succ 0\quad\mbox{and}\quad\widehat{V}_{0}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V_{0}\succ 0\quad\mbox{on }M\,.$
Notice that we are not assuming that, globally, $V$ and $V_{0}$ are non-negative. This will be done in Assumption 8.3, which is, hence slightly stronger than Assumption 8.1. We also recall that, if $a$ is a multiplication operator, then we use also the notation $\widehat{a}$ for ${\mathcal{R}_{\infty}}(a)$ (see Equation (5.6) for a justification of this notation). We also recall (see Definition 2.13) that the explicit forms of the last two assumptions amount to the following:
@@
    $\widehat{V}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\widehat{V}(x)$ is invertible (i.e., $V_{0}(x)>0$ in $M_{n}(\mathbb{C})$).
@@
    $\widehat{V}_{0}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\widehat{V}_{0}(x)$ is invertible (i.e., $V_{0}(x)\in(0,\infty)$).
@@ 8.2. The Moore-Penrose pseudoinverse of $\boldsymbol{\Xi}$
Let $\mathcal{N}$ be the kernel of $\boldsymbol{\Xi}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$. The classical elliptic regularity of $\boldsymbol{\Xi}$ implies that $\mathcal{N}$ will consist of smooth sections. Assume that $\mathcal{N}\subset L^{2}(M;TM\oplus\mathbb{C})$ and that it is finite dimensional and let $p_{\mathcal{N}}$ be the $L^{2}$–orthogonal projection onto $\mathcal{N}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$ in Definition 6.11. Assume also that there exists a pseudodifferential operator
@@
Let $M$ be a manifold with straight cylindrical ends and assume that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ (Assumption 8.1). Then the following properties hold.
@@
Assumption 8.1 implies that ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$ satisfies the assumptions of Theorem 7.14, which then yields right away the first point. The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is elliptic by Proposition 6.13. Since ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ is invertible, Theorem 6.15 implies that $\boldsymbol{\Xi}$ is Fredholm. The last point follows from the Beals-type result proved in [87, Theorem 7.4] (this result can also be found in [83]). ∎
@@
Let $U\in H^{1}(M;TM)\oplus L^{2}(M)$ be such that $\boldsymbol{\Xi}U=0$ on $M$. Then $U\in H^{2}(M;TM)\oplus H^{1}(M)$, by elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]). Let us notice that the fact that $V$ and $V_{0}$ satisfy the non-vanishing assumption on $M$ (Assumption 8.3) and that $M$ is connected implies that $V_{0}\not\equiv 0$ and $V\succ 0$ on $M$. The assumption $\boldsymbol{\Xi}U=0$ then allows us to use Corollary 2.15 to conclude that $U=0$. Thus $\boldsymbol{\Xi}$ is invertible between the indicated spaces, and hence also as a pseudodifferential operator. The fact that its inverse belongs to the set $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ follows from a “Beals-type theorem” (Theorem 7.4 in [87]). The last stated isomorphism follows from the mapping properties of operators in the spaces $\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM)$ and in $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM)$, as in the last corollaries. This completes the proof. ∎
The invertibility of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ in the case $M$ closed is obtained when both $V$ and $V_{0}$ do not vanish identically on $M$. Moreover, the condition that $V$ does not vanish identically on $M$ can be replaced with the condition that $M$ does not have non-vanishing Killing vector fields. These results were proved in [89, Theorem 6.1(4)]. In relation to this result, the reader might have noticed that we are not considering the case when $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ has straight cylindrical ends but does not have non-vanishing Killing vector fields. In fact, this condition is not very useful for manifolds with straight cylindrical ends because, even if $M$ does not have non-vanishing Killing vector fields, its “end” $M^{\prime}\times\mathbb{R}$ will always have non-vanishing Killing vector fields, for instance $\partial_{t}$ (the infinitesimal generator of the translations along $\mathbb{R}$). (See Subsection 12.2 for more on the case $M$ closed.)
@@ 8.4. Jump relations
For the Stokes operator $\boldsymbol{\Xi}_{0,0}$, see also[49], [82, Lemma 3.1], [90, (6.1), (6.2)], [121, Theorem 3.1], [122], [158, Lemma 1.3]. Note that $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\frac{1}{2}+\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$. Therefore $\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\boldsymbol{S}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$. In addition, we obtain the following formula.
@@ 9. Calculations on a cylinder
We shall need definitions and results for the limit operators ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ that are analogous to those in the previous two sections. To this end, in this section we consider the particular case when our manifold $M$ with straight cylindrical ends is actually a cylinder $M=\mathfrak{S}\times\mathbb{R}$ with a product metric. We do not assume $\mathfrak{S}$ to be connected, though. Then, as in Section 7, we let $\Omega=\Omega^{\prime}\times\mathbb{R}\subset M$ be a product type open subset, and hence with straight cylindrical ends, where $\Omega^{\prime}\subset\mathfrak{S}$. Let $\Gamma^{\prime}:=\partial\Omega^{\prime}$. In this section, in addition to $V$ and $V\geq 0$ smooth, we assume that they are translation invariant, and hence $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$. Also, in this section, we use Assumption 8.1, that is that $\widehat{V}\succ 0$ and that $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (that is, they have points where they are strictly positive in every connected component of the space of definition, see Definition 2.13). Also, recall that $\Omega$ is on one side of its boundary (Assumption 2.7), and hence $\Omega^{\prime}$ is also on one side of its boundary.
@@ 9.1. Layer potentials and jump relations for the indicial operators
Assume that $\widehat{V}\succ 0$ and that $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (Assumption 8.1). Let $\boldsymbol{h}\in H^{s}(\Gamma^{\prime};TM)$, $s\in\mathbb{R}$. The indicial layer potentials operators associated to $\widehat{\boldsymbol{\Xi}}(\tau)$ are defined by
@@
Let $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$, $\Gamma^{\prime}:=\partial\Omega^{\prime}$, $\widehat{\mathfrak{f}}:=\frac{\widehat{V}_{0}+1}{2\widehat{V}_{0}+1}$ $\widehat{\mathfrak{g}}:=\frac{1}{2\widehat{V}_{0}+1}$. Let $\boldsymbol{K}:=\boldsymbol{P}_{0}\in\Psi^{0}(\Gamma^{\prime};TM)$ be as in Equation (8.6). Then
@@
    $\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\,=\,\widehat{\mathcal{A}}_{0}(\tau)\boldsymbol{h}^{\prime}$, and $\widehat{\mathcal{A}}_{0}(\tau)$ is elliptic with self-adjoint symbol $\sigma_{-1}(\mathcal{A}_{0}(\tau);\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\widehat{\mathfrak{f}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\widehat{\mathfrak{f}}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$.
@@
    $\mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}=\big(\mp\frac{\widehat{\mathfrak{g}}}{2}\boldsymbol{\nu}^{\sharp}+\widehat{\mathcal{C}}_{0}(\tau)\big)\boldsymbol{h}^{\prime}$.
@@ 9.2. Invertibility of $\boldsymbol{S}$ on a cylinder
  • $V$ and $V\geq 0$ are smooth, translation invariant, and hence $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
  • $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1, that is, $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$).
@@
    $\widehat{V}\succ 0$ on $\mathfrak{S}$ and
@@
    $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
@@
Let us show now that $\ker\widehat{\boldsymbol{S}\,}(0)=0$. Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Our assumptions imply that $\widehat{\boldsymbol{\Xi}}(0)$ is invertible (Proposition 7.11), and hence the indicial single layer potential $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{0}(\boldsymbol{h}^{\prime})$ is defined (see Definition 9.2). Since $\widehat{\boldsymbol{S}\,}(0)$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by classical elliptic regularity. Proposition 9.5 for $s=1/2$ then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}{\,}^{\prime}$ satisfy
@@
We also know that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{\pm}^{\prime}$ by Proposition 8.10. Proposition 9.6 gives then that $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Therefore $\boldsymbol{u}=0$ in $\Omega_{-}^{\prime}$ and in $\Omega_{+}^{\prime}$, by Corollary 7.7. The same corollary gives that $p$ is constant on each connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. Because $\widehat{V}_{0}$ is not identically equal to zero on any connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$, we obtain that, in fact, $p=0$. We have thus obtained that $U=0$ and hence $\widehat{\boldsymbol{S}\,}(0)$ is injective.
@@
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $V\succ 0$ and $V_{0}\succ 0$ on $M$ (equivalently, that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$).
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $V$ and $V_{0}$ are smooth, translation invariant, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
@@
The proof is exactly the same as that of Proposition 9.7, except that we are using Corollary 7.9 (which only requires $\widehat{V}$ and $\widehat{V}_{0}$ to be non-negative) instead of Corollary 7.7. ∎
@@
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\widehat{V}\succ 0$ on $\mathfrak{S}$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. Then $\boldsymbol{S}$ is invertible, $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, and hence $\boldsymbol{S}:H^{s}(\Gamma;TM)\to H^{s+1}(\Gamma;TM)$ is an isomorphism for all $s\in\mathbb{R}$.